# Splitting of uniform bundles on quadrics

### v1

@@ 3.3. The case $l$ is odd
Appendix A Classification of primitive approximate solutions
@@
Let $S_{0}(T,X_{1},X_{2})$ be a symmetrical divisor of $P(T,X_{1})-R(X_{1},X_{2})$ such that for a $2(n+1)$-th primitive unit root $y_{0}$, we have $S_{0}(0,1,y_{0})=0$. Since $y_{0}$ is a simple root of $P(0,1)-R(1,z)$, $y_{0}$ is also a simple root of $S_{0}(0,1,z)$. Therefore, by the implicit function theorem, there is a germ of holomorphic function $y(x)$ at a neighborhood of $x=0$ satisfying
@@
We now show by induction that for $m=1,\dots,2n-1$, we have $y^{(m)}(0)=p_{2n-m}=0$. As $S_{0}$ is symmetrical in $X_{1}$ and $X_{2}$, we have
@@
By taking the derivatives of (1) and (2) at $x=0$ and note that $p_{2n}$ is 00, we obtain
@@
From $R(y,1)(y^{2}-1)=y^{2n+2}-1$, we have $R^{\prime}(y,1)(y^{2}-1)+R(y,1)\cdot(2y)=(2n+2)y^{2n+1}$. As $R(y_{0},1)$ is 00 and $y_{0}$ is not ±$\pm 1$, $R^{\prime}(y_{0},1)$ is $\frac{(2n+2)y_{0}^{2n+1}}{y_{0}^{2}-1}(\neq 0)$. From (1’) and (2’), we get $p_{2n-1}(1+y_{0})(y_{0}^{2n-1}-1)=0$. Since $y_{0}$ is primitive, $y_{0}^{2n-1}-1$ and $1+y_{0}$ are not zero. We have $p_{2n-1}=0$. By (1’) and note $R^{\prime}(y_{0},1)\neq 0$, $y^{\prime}(0)$ is 00.
@@
Let $S_{+}(T,X_{1},X_{2})$ be a symmetrical divisor of $P(T,X_{1})-R(X_{1},X_{2})$ such that for a $2n$-th primitive unit root $y_{0}$, we have $S_{+}(0,1,y_{0})=0$. Note the equation $\Sigma_{n}(p_{2n-2}T^{2},X_{1}^{2})-R(X_{1},X_{2})=(p_{2n-2}T^{2}-X_{2}^{2})\Sigma_{n-1}(p_{2n-2}T^{2},X_{1}^{2},X_{2}^{2})$ (see [2, Section 7.2] for example). Let $S_{-}(T,X_{1},X_{2})$ be $\Sigma_{n-1}(p_{2n-2}T^{2},X_{1}^{2},X_{2}^{2})$, then $S_{-}(0,1,y_{0})$ is also 00. Denote $P(T,X_{1})$ by $P_{+}(T,X_{1})$ and denote $\Sigma_{n}(p_{2n-2}T^{2},X_{1}^{2})$ by $P_{-}(T,X_{1})$, we are going to show that $P_{+}$ equals $P_{-}$. Let $y_{\pm}(x)$ be germs of holomorphic functions satisfying
@@
As $S_{\pm}$ is symmetrical in $X_{1}$ and $X_{2}$, we have equations:
@@
Multiplying ($2^{\prime}_{\pm}$) by $y_{0}^{2}$ and using the relation (A), one has
@@
Since $y_{0}$ is primitive, $\left|\begin{array}[]{ccc}y_{0}(1+y_{0})&-R^{\prime}(y_{0},1)\\ 1+y_{0}&-R^{\prime}(y_{0},1)\end{array}\right|=(1-y_{0})(1+y_{0})R^{\prime}(y_{0},1)$ is not 00. Then one obtains $p_{2n-1}^{\pm}=y_{\pm}^{\prime}(0)=0$.
@@
where $T^{i}_{\pm}$ are the remaining terms satisfying $T^{1}_{-}=T^{1}_{+}$ and $T^{2}_{-}=T^{2}_{+}$. For the case $m=2$, by construction, we automatically have $p_{2n-2}^{-}=p_{2n-2}=p_{2n-2}^{+}$. Then one obtains $y_{+}^{(2)}(0)=y_{-}^{(2)}(0)$ from ($1^{m}_{\pm}$). Suppose now $m$ is at least $3$, multiplying ($2^{m}_{\pm}$) by $y_{0}^{2}$ and using (A), we have

### v2

@@ 3.3. The case $l$ is odd
4. Splitting type of unsplit uniform bundle of minimal rank
There are some restrictions on the splitting types of unsplit uniform bundles of minimal rank. The following theorem generalizes [1, Corollary 4.7] to generalized Grassmannians associated with short roots.
Theorem 4.1.
Assume $E$ is an unsplit uniform bundle on a generalized Grassmannian $X$ whose rank is $\mu(X)+1$. If the splitting type of $E$ is $(a_{1},a_{2},\dots,a_{r})$ $(a_{1}\geq a_{2}\geq\cdots\geq a_{r})$, then $max\{a_{i}-a_{i+1}|1\leq i\leq r-1\}$ is $1$.
We now prove Theorem 4.1.
Proof.
Let $L$ be a line in $X$. We denote the inclusion morphism by $f_{L}:L(\simeq\mathbb{P}^{1})\rightarrow X$. As $X$ is a homogeneous variety, the tangent bundle $T_{X}$ is globally generated. Then $f_{L}^{*}(T_{X})$ is also globally generated and hence $H^{1}(L,f_{L}^{*}(T_{X}))$ vanishes. So Mor$(\mathbb{P}^{1},X)$ is smooth at $[f_{L}]$.
If $max\{a_{i}-a_{i+1}|1\leq i\leq r-1\}$ is 00, by [10, Theorem 1.2], $E$ is trivial. If $max\{a_{i}-a_{i+1}|1\leq i\leq r-1\}$ is at least $2$, there would exist a number $j<r$ such that $a_{j}-a_{j+1}\geq 2$. By [11, Proposition 3.1], there would be a subbundle $W$ of $E$ satisfying the following two properties:
  • $W$ is a uniform bundle of splitting type $(a_{1},\dots,a_{j})$.
  • the quotient $U\triangleq E/W$ is a uniform bundle of splitting type $(a_{j+1},\dots,a_{r})$.
As rk$(U)$ and rk$(W)$ are at most $\mu(X)$, both $U$ and $W$ split. Since $\text{Ext}^{1}(U,W)$ vanishes, the exact sequence $0\rightarrow W\rightarrow E\rightarrow U\rightarrow 0$ would split. So $E(\simeq U\oplus W)$ would be a direct sum of line bundles. ∎
In [3, Page 29, Conjecture], Ellia proposes a conjecture that every uniform bundle on $\mathbb{P}^{n}$ of splitting type
$(\underbrace{a_{1},\dots,a_{1}}_{l_{1}},\underbrace{a_{2},\dots,a_{2}}_{l_{2}},\dots,\underbrace{a_{k},\dots,a_{k}}_{l_{k}})$
with $a_{i}>a_{i+1}$ and $l_{i}\leq n-1$ for any $1\leq i\leq k$ necessarily splits. Using the same argument as in Theorem 4.1, we can reduce this conjecture to the case $l_{i}\leq n-1$ and $a_{i}-a_{i+1}=1$ for any $11\leq i\leq k-$ 1. In particular, the following Theorem 4.2 partially answers Ellia’s conjecture in a more general setting.
Theorem 4.2.
Given a generalized Grassmannian $X$. Let $E$ be a uniform bundle on $X$. Assume that the splitting type of $E$ is $(\underbrace{a_{1},\dots,a_{1}}_{l_{1}},\underbrace{a_{2},\dots,a_{2}}_{l_{2}},\dots,\underbrace{a_{k},\dots,a_{k}}_{l_{k}})$ with $a_{i}>a_{i+1}$. If for any $11\leq i\leq k-$ 1, $a_{i}-a_{i+1}$ is at least $2$, then $E$ splits.
Proof.
We prove by induction on $k$. The case $k=1$ is obviously true. Suppose the case $k<n$ is true. We now prove the case $k=n$.
Following the same strategy in the proof of Theorem 4.1, by [11, Proposition 3.1], there exists a subbundle $W$ of $E$ satisfying the following two properties:
  • $W$ is a uniform bundle of splitting type $(\underbrace{a_{1},\dots,a_{1}}_{l_{1}})$.
  • the quotient $U\triangleq E/W$ is a uniform bundle of splitting type $(\underbrace{a_{2},\dots,a_{2}}_{l_{2}},\dots,\underbrace{a_{n},\dots,a_{n}}_{l_{n}})$
Then $W$ is isomorphic to $\mathcal{O}_{X}^{\oplus l_{1}}(a_{1})$ where $\mathcal{O}_{X}(1)$ is the ample generator of Pic$(X)$. By induction, the bundle $U$ splits. Since $\text{Ext}^{1}(U,W)$ vanishes, the exact sequence $0\rightarrow W\rightarrow E\rightarrow U\rightarrow 0$ splits. So $E(\simeq U\oplus W)$ is a direct sum of line bundles.
Appendix A Additional details for Proposition 2.1
In this appendix, we use notations as in the proof of Proposition 2.1 and provide more details by sketching the proof of [14, Page 204, Case (ii)] and [13, Lemma 3.3 (ii) and Case 1 of (iii)].
Suppose that there exists a non-constant morphism $f:\mathbb{Q}^{2n+1}\rightarrow\mathbb{G}(l,2n+1)$. Let $g(t)$ and $h(t)$ be the polynomials in Proposition 2.1. We set
$g(t)=1-c_{1}t+c_{2}t^{2}+\cdots+(-1)^{l+1}c_{l+1}t^{l+1},$
$h(t)=1+d_{1}t+d_{2}t^{2}+\cdots+d_{2n+1-l}t^{2n+1-l}.$
Let $a$ be $2n+2]{c_{l+1}d_{2n+1-l}}$. Then, we have the equation (see Page 3, the equation (2.1))
$g(t)h(t)=1+(-1)^{l+1}c_{l+1}d_{2n+1-l}t^{2n+2}=1+(-1)^{l+1}a^{2n+2}t^{2n+2}.$ (A.1)
First, we use the same method of [13, Lemma 3.3] to prove that $a$ is a rational number.
Proposition A.1.
We have the identities $t^{l+1}g(\frac{1}{at})=(-1)^{l+1}g(\frac{t}{a})$ and $t^{2n+1-l}h(\frac{1}{at})=h(\frac{t}{a})$.
Proof.
Set $g(t)=(1-\alpha_{1}at)(1-\alpha_{2}at)\cdots(1-\alpha_{l+1}at)$. By formula (A.1), we have
$|\alpha_{i}|=1,~{}\alpha_{i}\neq\alpha_{j}\text{ if }i\neq j,~{}\alpha_{i}^{-1}=\overline{\alpha_{i}}\in\{\alpha_{1},\alpha_{2},\dots,\alpha_{l+1}\}.$ (A.2)
Furthermore, $\alpha_{1}\cdots\alpha_{l+1}a^{l+1}$ equals $c_{l+1}$. Since both $a$ and $c_{l+1}$ are positive real numbers, we must have $\alpha_{1}\cdots\alpha_{l+1}=1$. Combining with (A.2), we get
$t^{l+1}g(\frac{1}{at})$ $=(t-\alpha_{1})(t-\alpha_{2})\cdots(t-\alpha_{l+1})=(\alpha_{1}^{-1}t-1)(\alpha_{2}^{-1}t-1)\cdots(\alpha_{l+1}^{-1}t-1)\alpha_{1}\cdots\alpha_{l+1}$
$=(\alpha_{1}t-1)(\alpha_{2}t-1)\cdots(\alpha_{l+1}t-1)=(-1)^{l+1}g(\frac{t}{a}).$
Similarly, we set $h(t)=(1+\alpha_{1}^{\prime}at)(1+\alpha_{2}^{\prime}at)\cdots(1+\alpha_{2n+1-l}^{\prime}at)$ and apply similar arguments to obtain $t^{2n+1-l}h(\frac{1}{at})=h(\frac{t}{a})$. ∎
Corollary A.2.
The number $a$ is $\frac{c_{m+1}}{c_{m}}$, where $m$ is $\frac{l}{2}$.
Proof.
The equation $t^{l+1}g(\frac{1}{at})=(-1)^{l+1}g(\frac{t}{a})$ shows that $c_{i}a^{-i}$ equals $c_{l+1-i}a^{-l-1+i}$ for $1\leq i\leq l$. We get the desired conclusion by taking $i$ to be $\frac{l}{2}$. ∎
Once one proves that $a$ is rational, as in the proof of Proposition 2.1, both $a,C_{i}$ and $D_{j}$ are positive integers, we claim that
Claim: If $a,C_{i}$ and $D_{j}$ are all positive integers, we can apply the same proof of [14, Case (ii)] to get a contradiction.
Now we sketch the proof of [14, Page 204, Case (ii)]. Let $G(t)$ be $g(\frac{t}{a})(=1-C_{1}t+C_{2}t^{2}+\cdots+(-1)^{l+1}C_{l+1}t^{l+1})$ and $H(t)$ be $h(\frac{t}{a})(=1+D_{1}t+D_{2}t^{2}+\cdots+D_{2n+1-l}t^{2n+1-l})$. From formula (A.1), we have $H(t)G(t)=1+(-1)^{l+1}t^{2n+2}$.
Lemma A.3.
Suppose that there exists a non-constant morphism $f$ from $\mathbb{Q}^{2n+1}$ to $\mathbb{G}(l,2n+1)$. Then we have the equations $n=l$, $C_{1}=C_{2}=\cdots=C_{l}=D_{1}=D_{2}=\cdots=D_{2n-l}=2$ and $C_{l+1}=D_{2n+1-l}=1$.
Proof.
Since $\mathbb{G}(l,2n+1)$ is isomorphic to $\mathbb{G}(2n+1-l,2n+1)$, we may assume the inequality $n\geq l$. From the equation $H(t)G(t)=1+(-1)^{l+1}t^{2n+2}$ and the fact that $C_{i},D_{j}$ are positive integers, the first part of Tango’s proof allows us to conclude that $H(1)$ equals $22n+$ 2 and $D_{1}$ is bigger than $1$.
Next we prove that $D_{1},D_{2},\dots,D_{2n-l}$ are not less than $2$. Suppose that there exists a positive integer $k(\leq 2n-l)$ such that $D_{k}$ equals $1$. Let $r$ be $\min\{k|D_{k}=1\}$. Then both $D_{1}$ and $D_{r-1}$ are bigger than $1$. From $H(t)=t^{2n+1-l}H(\frac{1}{t})$ (Proposition A.1), we obtain $D_{2n+1-l}=1$ and $D_{2n+1-i}=D_{i}$ for $1\leq i\leq 2n-l$. In particular, $D_{2n+1-r}$ is $D_{r}$. By the definition of $r$, we have $2r\leq 2n+1-l$. Let $Q$ be the universal quotient bundle of $\mathbb{G}(l,2n+1)$. By Pieri’s formula, the class
$f^{*}\omega_{r,r,0,\dots,0}$ $=f^{*}(\omega_{r,0,\dots,0}^{2}-\omega_{r+1,0,\dots,0}\omega_{r-1,0,\dots,0})$
$=f^{*}(c_{r}(Q))^{2}-f^{*}(c_{r+1}(Q))f^{*}(c_{r-1}(Q))=(D_{r}^{2}-D_{r+1}D_{r-1})a^{2r}H^{2r}$
is the pullback of a Schubert cycle, which is numerically non-negative. (For the definition of the Schubert cycle $\omega_{a_{0},\dots,a_{l}}$, we refer to [14]). So $D_{r}^{2}-D_{r+1}D_{r-1}\geq 0$. From the inequalities $D_{r}^{2}-D_{r+1}D_{r-1}\leq 1-2<0$, we get a contradiction. Hence, we have $2n+2=H(1)=1+D_{1}+\cdots+D_{2n+1-l}\geq 1+2(2n-l)+1=2(2n+1-l)$. Combining with the assumption $n\geq l$, we must have $n=l$, $D_{1}=D_{2}=\cdots=D_{2n-l}=2$ and $D_{2n+1-l}=1$. Finally, from $H(t)G(t)=1+(-1)^{l+1}t^{2n+2}$, we also have $C_{1}=C_{2}=\cdots=C_{l}=2$ and $C_{l+1}=1$. ∎
Now we return to the proof of Claim. Since $l$ is even and $(l,2n+1)$ is not $(2,5)$, we have $l\geq 4$ and $D_{1}=D_{2}=D_{3}=2$. Similar as above, we have $f^{*}\omega_{2,2,0,\dots,0}=f^{*}(\omega_{2,0,\dots,0}^{2}-\omega_{3,0,\dots,0}\omega_{1,0,\dots,0})=(D_{2}^{2}-D_{3}D_{1})a^{4}H^{4}=0$. By Pieri’s formula again, it shows that
$0=f^{*}(\omega_{2,2,0,\dots,0}\omega_{2n-1-l,0,\dots,0})=f^{*}\omega_{2n+1-l,2,0,\dots,0}+f^{*}\omega_{2n-l,2,1,0,\dots,0}+f^{*}\omega_{2n-l-1,2,2,0,\dots,0}.$
Since all the classes in the right hand side is numerically non-negative, we have $f^{*}\omega_{2n+1-l,2,0,\dots,0}=0$. On the other hand, we have
$f^{*}\omega_{2n+1-l,2,0,\dots,0}=f^{*}\omega_{2n-1-l,0,\dots,0}f^{*}\omega_{2,0,\dots,0}$
$=$ $f^{*}(c_{2n+1-l}(Q))f^{*}(c_{2}(Q))=D_{2n+1-l}D_{2}a^{2n+3-l}H^{2n+3-l}\neq 0,$
which is a contradiction.
Appendix B Classification of primitive approximate solutions
@@
Let $S_{0}(T,X_{1},X_{2})$ be a symmetric divisor of $P(T,X_{1})-R(X_{1},X_{2})$ such that for a $2(n+1)$-th primitive unit root $y_{0}$, we have $S_{0}(0,1,y_{0})=0$. Since $y_{0}$ is a simple root of $P(0,1)-R(1,z)$, $y_{0}$ is also a simple root of $S_{0}(0,1,z)$. Therefore, by the implicit function theorem, there is a germ of holomorphic function $y(x)$ in a neighborhood of $x=0$ satisfying
@@
We now show by induction that for $m=1,\dots,2n-1$, we have $y^{(m)}(0)=p_{2n-m}=0$. As $S_{0}$ is symmetric in $X_{1}$ and $X_{2}$, we have
@@
By taking the derivatives of (1) and (2) at $x=0$ and noting that $p_{2n}$ is 00, we obtain
@@
From $R(y,1)(y^{2}-1)=y^{2n+2}-1$, we have $R^{\prime}(y,1)(y^{2}-1)+R(y,1)\cdot(2y)=(2n+2)y^{2n+1}$. As $R(y_{0},1)$ is 00 and $y_{0}$ is not ±$\pm 1$, $R^{\prime}(y_{0},1)$ is $\frac{(2n+2)y_{0}^{2n+1}}{y_{0}^{2}-1}(\neq 0)$. From (1’) and (2’), we get $p_{2n-1}(1+y_{0})(y_{0}^{2n-1}-1)=0$. Since $y_{0}$ is primitive, $y_{0}^{2n-1}-1$ and $1+y_{0}$ are not zero. We have $p_{2n-1}=0$. By (1’) and noting that $R^{\prime}(y_{0},1)\neq 0$, $y^{\prime}(0)$ is 00.
@@
Let $S_{+}(T,X_{1},X_{2})$ be a symmetric divisor of $P(T,X_{1})-R(X_{1},X_{2})$ such that for a $2n$-th primitive unit root $y_{0}$, we have $S_{+}(0,1,y_{0})=0$. Note the equation $\Sigma_{n}(p_{2n-2}T^{2},X_{1}^{2})-R(X_{1},X_{2})=(p_{2n-2}T^{2}-X_{2}^{2})\Sigma_{n-1}(p_{2n-2}T^{2},X_{1}^{2},X_{2}^{2})$ (see [2, Section 7.2] for example). Let $S_{-}(T,X_{1},X_{2})$ be $\Sigma_{n-1}(p_{2n-2}T^{2},X_{1}^{2},X_{2}^{2})$, then $S_{-}(0,1,y_{0})$ is also 00. Denote $P(T,X_{1})$ by $P_{+}(T,X_{1})$ and denote $\Sigma_{n}(p_{2n-2}T^{2},X_{1}^{2})$ by $P_{-}(T,X_{1})$. We are going to show that $P_{+}$ equals $P_{-}$. Let $y_{\pm}(x)$ be germs of holomorphic functions satisfying
@@
As $S_{\pm}$ is symmetric in $X_{1}$ and $X_{2}$, we have equations:
@@
Multiplying ($2^{\prime}_{\pm}$) by $y_{0}^{2}$ and using the relation (B), one has
@@
Since $y_{0}$ is primitive, the number $\left|\begin{array}[]{ccc}y_{0}(1+y_{0})&-R^{\prime}(y_{0},1)\\ 1+y_{0}&-R^{\prime}(y_{0},1)\end{array}\right|=(1-y_{0})(1+y_{0})R^{\prime}(y_{0},1)$ is not 00. Then one obtains $p_{2n-1}^{\pm}=y_{\pm}^{\prime}(0)=0$.
@@
where $T^{i}_{\pm}$ are the remaining terms satisfying $T^{1}_{-}=T^{1}_{+}$ and $T^{2}_{-}=T^{2}_{+}$. For the case $m=2$, by construction, we automatically have $p_{2n-2}^{-}=p_{2n-2}=p_{2n-2}^{+}$. Then one obtains $y_{+}^{(2)}(0)=y_{-}^{(2)}(0)$ from ($1^{m}_{\pm}$). Suppose now $m$ is at least $3$, multiplying ($2^{m}_{\pm}$) by $y_{0}^{2}$ and using (B), we have